PLEASEHELP AND SHOW WORK.IF THERE IS A WAY ON THE TI84 PLEASE SHOW
Test the claim that the proportion of people who own cats is smaller than 10% at the 0.005 significance level The null and alternative hypothesis would be: Ho:p 0.1 H0 : p> 0.1 Ho:p=0.1 H0μ 0.1 Ha:p> 0.1 Hai p > 0.1 Ha: p < 0.1 HUP 0.1 Ho: μ-0.1 Ho: μ > 0.1 Ha : μ 0.1 Ha : μ < 0.1 The test is: right-tailed two-tailed left-tailed Based on a sample of 100 people, 6% owned cats The p-value is: (to 2 decimals) Based on this we: OReject the null hypothesis OFail to reject the null hypothesis
Answer #1

Answer

claim is that the proportion of people who own cats is less than 10%

So, for null hypothesis we assume that the proportion is equal to 10% and for alternate hypothesis, we assume that the proportion is less than 10%

we can write it as

H_o: p ge 0.10

H_a: p < 0.10

So, option B is correct answer

Since the alternate hypothesis include less than symbol, so it is left tailed hypothesis.

Sample size is given as n = 100 and sample proportion = 6% / 6/100 = 0.06

population proportion = 10% = 10/100 = 0.10

Formula for test statistic is given as

z = (hat{p}-p_o)/sqrt{(p_o*(1-p_o))/n}

setting the given values, we get

z = (0.06-0.10)/sqrt{(0.10*(1-0.10))/100} = -0.04/0.03 = -1.33

Now, we have to use z distribution to get the required p value. Look at -1.0 in left most column and 0.33 in the top most row.

we get

P(z <-1.33) = 1-P(z < 1.33) = 1-0.9082 = 0.0918

rounding to 2 decimals, we get p value = 0.09

At 0.005 level of significance, p value is greater than significance level, so result is insignificant.

We failed to reject the null hypothesis.